# Segmenting Action-Value Functions Over Time-Scales in SARSA using TD($\Delta$)

### v1

@@ 4 Methodology
The TD($\Delta$) framework enhances conventional temporal difference learning by partitioning action-value functions across several discount factors $\gamma_{0},\gamma_{1},...,\gamma_{z}$. This decomposition enables the action-value function to be acquired as a summation of delta estimators, each representing the disparity between action-value functions at varying discount factors. The principal advantage of this method is the capacity to regulate variance and bias in the learning process by concentrating on smaller, more manageable time-scales.
@@ 4.1 TD($\Delta$) Framework
To calculate $W_{z}$ for SARSA as outlined in the study, it is essential to comprehend how delta estimators ($W_{z}$) approximate the discrepancies between action-value functions at consecutive discount factors. Let us delineate this approach for SARSA in a stepwise manner.
Delta Estimators ($W_{z}$) and Action-Value Functions ($Q(s,a)$): Delta estimators, $W_{z}$, represent the difference between value functions with successive discount factors:
$W_{z}=Q_{\gamma_{z}}\mathit{-}Q_{\gamma_{z-1}}$ (5)
In this context, $Q_{\gamma_{z}}$ denotes the action-value function with discount factor $\gamma_{z}$, but $Q_{\gamma_{z-1}}$ signifies the action-value function with the prior discount factor, $\gamma_{z-1}$.
@@ 4.2 Single-Step TD ( SARSA($\Delta$) )
This section begins with a description of the delta estimators ($W_{z}$) for the action-value functions $Q$ utilized in SARSA. The extension of TD($\Delta$) to SARSA is denoted as SARSA($\Delta$). SARSA($\Delta$) aims to obtain the action-value function $Q(s,a)$, representing the expected return of adhering to policy $\pi$ from state $s$, performing action $a$, and subsequently continuing to follow $\pi$. The aim is to improve the overall reward through the optimization of the action-value function across different time-scales. The core principle in SARSA($\Delta$) involves dividing the action-value function $Q(s,a)$ into a series of delta components $W_{z}(s,a)$, with each component associated with a specific discount factor $\gamma_{z}$. The interrelation among these delta components (i.e., delta function) is articulated as follows:
$W_{z}(s,a):=Q_{\gamma_{z}}(s,a)\mathit{-}Q_{\gamma_{z-1}}(s,a)$ (6)
where $\gamma_{0},\gamma_{1},...,\gamma_{z}$ represent the discount factors across various time-scales, and define $W_{0}(s,a):=Q_{\gamma_{0}}(s,a)$. The action-value function $Q_{\gamma_{z}}(s,a)$ is defined as the cumulative sum of all W-components up to z:
$Q_{\gamma_{z}}(s,a)=\sum_{j=0}^{z}W_{j}(s,a)$ (7)
@@
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t}+\gamma Q(s_{t+1},a_{t+1})-Q(s_{t},a_{t})\right]$ (8)
Here, $r_{t}$ denotes the reward obtained from executing action $a_{t}$ in state $s_{t}$, while $(s_{t+1},a_{t+1})$ represents the subsequent state-action pair. The SARSA update rule is applied individually to each delta component $W_{z}(s,a)$. The update for each component employs the same framework as the traditional SARSA method, yet it is adjusted to incorporate the delta function at each specific time-scale. We can express the update Eq.8 for single-step TD SARSA ($\Delta$) with multiple time-scales by applying the concept of separating the update into multiple discount factors[8] as follows:
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}W_{z}(s_{t+1},a_{t+1})\right]$ (9)
where $Q_{\gamma_{z}}(s_{t},a_{t})$ represents the action-value function with a discount factor $\gamma_{z}$. $W_{z}(s_{t},a_{t})$ denotes the delta function for SARSA ($\Delta$), while $\gamma_{z}$ represents the discount factor for time-scale z, while $Q_{\gamma_{z}}(s_{t+1},a_{t+1})$ is the SARSA update for action-value associated with the subsequent state-action pair. The single-step TD SARSA ($\Delta$) mirrors the original TD update, but it is applied to action-value functions Q rather than state-value functions V. The single-step Bellman Eq. for the Q-value function in standard SARSA is as follows:
$Q_{\gamma_{z}}(s_{t},a_{t})=\mathbb{E}\left[r_{t}+\gamma_{z}Q_{\gamma_{z}}(s_{t+1},a_{t+1})\right]$ (10)
This denotes the expected value of the reward acquired at time t, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{t+1},a_{t+1})$. Currently, rather than utilizing a singular discount factor $\gamma$, we are employing a sequence of discount factors $\gamma_{0},\gamma_{1},.....,\gamma_{z}$. The delta function $W_{z}(s_{t},a_{t})$ is defined as the difference between action-value functions associated with successive discount factors from Eq. 6:
$W_{z}(s_{t},a_{t})=Q_{\gamma_{z}}(s_{t},a_{t})-Q_{\gamma_{z-1}}(s_{t},a_{t})$ (11)
To expand the delta function $W_{z}$ utilizing its definition, we replace the Bellman Eq. for both $Q_{\gamma_{z}}$ and $Q_{\gamma_{z-1}}$, thereby expressing $W_{z}$ in relation to these action-value functions. Commence by composing:
@@
Since the reward terms $r_{t}$ are included in both formulations, simplify by eliminating them.
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}Q_{\gamma_{z}}(s_{t+1},a_{t+1})-\gamma_{z-1}Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
@@
$Q_{\gamma_{z}}(s_{t+1},a_{t+1})=W_{z}(s_{t+1},a_{t+1})+Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$
Substitute the definition of $W_{z}$ into the Eq. below.
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}(W_{z}(s_{t+1},a_{t+1})+Q_{\gamma_{z-1}}(s_{t+1},a_{t+1}))-\gamma_{z-1}Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
@@
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\gamma_{z}W_{z}(s_{t+1},a_{t+1})+(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})\right]$
Integrate the terms to derive the final update Eq. for $W_{z}(s_{t},a_{t})$.
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[(\gamma_{z}-\gamma_{z-1})Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}W_{z}(s_{t+1},a_{t+1})\right]$ (13)
This Eq. demonstrates that the update for $W_{z}$ integrates the disparity in discount factors $\gamma_{z}-\gamma_{z-1}$ with respect to the action-value function $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$ and the bootstrapping from the subsequent step $W_{z}(s_{t+1},a_{t+1})$. This denotes a Bellman Eq. for $W_{z}$, combining a decay factor $\gamma_{z}$ and the reward $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$. Consequently, it can be utilized to define the expected TD update for $W_{z}$. In this expression, $Q_{\gamma_{z-1}}(s_{t+1},a_{t+1})$ can be expressed as the summation of $W_{z}(s_{t+1},a_{t+1})$ for $z\leq Z-1$, indicating that the Bellman Eq. for $W_{z}$ is contingent upon the values of all delta functions $W_{z}$ for $z\leq Z-1$.
This approach treats the delta value function at each time-scale as an autonomous RL issue, with rewards obtained from the action-value function of the immediately preceding time-scale. Thus, for a target discounted action-value function $Q_{\gamma_{z}}(s,a)$, all delta components can be trained simultaneously through a TD update, utilizing the prior values of each estimator for bootstrapping. This process necessitates the assumption of a sequence of discount factors, denoted as $\gamma_{z}$, which includes both the minimum and maximum values, $\gamma_{0}$ and $\gamma_{z}$[8].
@@ 4.3 Multi-Step TD ( SARSA($\Delta$) )
Numerous studies indicate that multi-step TD methods typically exhibit greater efficiency compared to single-step TD methods [29]. In multi-step TD SARSA($\Delta$), rewards are aggregated over multiple steps instead of depending on a single future reward. This approach takes into account the differences between consecutive discount factors, $\gamma_{z}$ and $\gamma_{z-1}$, while utilizing the value function at state $s_{t+k_{z}}$ for bootstrapping. In SARSA($\Delta$), the update Eqs. are modified to aggregate rewards across a series of transitions prior to bootstrapping from either the Q-values or delta estimators. In multi-step SARSA($\Delta$), the agent aggregates rewards over K steps, leading to a corresponding modification in the update rule for each $W_{z}$.
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=1}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})\right]$ (14)
@@
In this case, $r_{t+j}$ represents the reward obtained at time step $t+j$. $a_{t+k_{z}}$ and the action is chosen greedily as $a=\arg\max_{a}Q(s_{t+k_{z}},a)$, where $k_{z}$ is the number of steps linked to the discount factor $\gamma_{z}$. This method generalizes Q-learning to multi-step TD learning by aggregating rewards across several stages and utilizing bootstrapping from both the current and prior time-scale action-value functions. In normal multi-step TD learning, the TD error is computed by aggregating rewards over several stages rather than relying on a single future step, followed by bootstrapping from the action-value at the last step. For example, utilizing a singular discount factor $\gamma$, the multi-step TD update equation is expressed as:
$Q(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k-1}\gamma^{j}r_{t+j}+\gamma^{k}Q(s_{t+k},a_{t+k})\right]$ (15)
@@
Herein, The initial segment of the Eq. aggregates the rewards across k steps, each diminished by $\gamma$. The second component derives from the action-value function at time step t+k. We expand this methodology to include various discount factors and delta decomposition, concentrating on representing the delta function $W_{z}(s_{t},a_{t})$ as the difference between action-value functions with discount factors $\gamma_{z}$ and $\gamma_{z-1}$. According to the definition of $W_{z}$, we obtain:
$W_{z}(s_{t},a_{t})=Q_{\gamma_{z}}(s_{t},a_{t})-Q_{\gamma_{z-1}}(s_{t},a_{t})$
@@
We now implement the multi-step variant of the Bellman Eq. for both $Q_{\gamma_{z}}(s_{t},a_{t})$ and $Q_{\gamma_{z-1}}(s_{t},a_{t})$. The multi-step Bellman Eq. for $Q_{\gamma_{z}}(s_{t},a_{t})$ is expressed as follows:
$Q_{\gamma_{z}}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}+\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
For $Q_{\gamma_{z-1}}(s_{t},a_{t})$, the multi-step Bellman Eq. is expressed as follows:
$Q_{\gamma_{z-1}}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}+\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
@@
$W_{z}(s_{t},a_{t})=\left[\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}+\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})\right]-\left[\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}+\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})\right]$
We will now broaden and simplify the expressions. The immediate reward terms $r_{t+j}$ are present in both Eqs. but are adjusted by distinct discount factors, $\gamma_{z}$ and $\gamma_{z-1}$. This enables us to express the disparity in rewards as:
$\sum_{j=0}^{k_{z}-1}\gamma_{z}^{j}r_{t+j}-\sum_{j=0}^{k_{z}-1}\gamma_{z-1}^{j}r_{t+j}=\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}$
@@
$\gamma_{z}^{k_{z}}Q_{\gamma_{z}}(s_{t+k_{z}},a_{t+k_{z}})-\gamma_{z-1}^{k_{z}}Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})=(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})$
@@
$W_{z}(s_{t},a_{t})=\mathbb{E}\left[\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})\right]$ (16)
This Eq. extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL. As a result, each $W_{z}$ obtains a share of the rewards from the environment up to time-step $k_{z}-1$. Furthermore, each $W_{z}$ utilizes its distinct action-value function in conjunction with the value from the prior time-scale. A variation of this algorithm, utilizing k-step bootstrapping as described in [29], is presented in Algorithm 1. Despite Algorithm 1 exhibiting quadratic complexity in relation to Z, it can be executed with linear complexity for substantial Z by preserving $\hat{Q}$ action-values at each time-scale $\gamma_{z}$.
@@
Inputs: Pick out the discount factors ($\gamma_{0},\gamma_{1},...,\gamma_{Z}$), bootstrapping steps ($k_{0},k_{1},...,k_{Z}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{Z}$).
Set the initial value of $W_{z}(s,a)=0$ for all states, actions, and scales z.
@@
     for t = 0, 1, 2,… do $\triangleright$ Loop for each time step
         Take action $a_{t}$, observe reward $r_{t}$ and next state $s_{t+1}$.
         Select action $a_{t+1}$ based on a policy.
         for z = 0, 1, …, Z do
              if z = 0 then
                  $G^{0}=\sum_{j=0}^{k_{0}-1}\gamma_{0}^{j}r_{t+j}+\gamma_{0}^{k_{0}}W_{0}(s_{t+k_{0}},a_{t+k_{0}})$
              $\triangleright$ Utilizing Eq.7, we substitute $Q_{\gamma_{z-1}}(s_{t+k_{z}},a_{t+k_{z}})$ by summing the W-components up to $W_{z-1}$ in Eq.4.3.
                  $G^{z}=\sum_{j=0}^{k_{z}-1}(\gamma_{z}^{j}-\gamma_{z-1}^{j})r_{t+j}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})\sum_{z=0}^{z-1}W_{z-1}(s_{t+k_{z}},a_{t+k_{z}})+\gamma_{z}^{k_{z}}W_{z}(s_{t+k_{z}},a_{t+k_{z}})$

### v2

@@ 4 Methodology
The TD($\Delta$) paradigm boosts regular TD learning by splitting action-value functions over several discount factors $\eta_{0},\eta_{1},...,\eta_{m}$. This decomposition facilitates the derivation of the action-value function as a sum of delta estimators, where each estimator indicates the distinction between action-value functions at various discount factors. Concentrating on smaller, more manageable time scales facilitates the regulation of learning variance and bias, which constitutes the primary benefit of this approach.
@@ 4.1 TD($\Delta$) Framework
To compute $D_{m}$ for SARSA as outlined in the study, it is essential to comprehend the mechanism by which delta estimators ($D_{m}$) approximate the discrepancies between action-value functions across successive discount factors. This approach for SARSA will be outlined in a stepwise manner.
Delta Estimators ($D_{m}$) and Action-Value Functions ($Q(s,a)$): Delta estimators, denoted as $D_{m}$, quantify the variation between action-value functions associated with consecutive discount factors:
$D_{m}=Q_{\eta_{m}}\mathit{-}Q_{\eta_{m-1}}$ (8)
In this context, $Q_{\eta_{m}}$ denotes the action-value function with discount factor $\eta_{m}$, but $Q_{\eta_{m-1}}$ signifies the action-value function with the prior discount factor, $\eta_{m-1}$.
@@ 4.2 Single-Step TD ( SARSA($\Delta$) )
This section initiates with an overview of the delta estimators ($D_{m}$) pertaining to the action-value functions $Q$ employed in SARSA. SARSA is extended to TD($\Delta$), which is referred to as SARSA($\Delta$). SARSA($\Delta$) seeks to derive the action-value function $Q(s,a)$, which indicates the anticipated return of following policy $\pi$ from state $s$, executing action $a$, and then persistently adhering to $\pi$ thereafter. The objective is to enhance the total reward by optimizing the action-value function over various time-scales. The fundamental concept in SARSA($\Delta$) is the segmentation of the action-value function $Q(s,a)$ into multiple delta components $D_{m}(s,a)$, where each component is linked to a distinct discount factor $\eta_{m}$. The relationship between these delta components (i.e., delta function) is expressed in the following manner:
$D_{m}(s,a):=Q_{\eta_{m}}(s,a)\mathit{-}Q_{\eta_{m-1}}(s,a)$ (9)
where $\eta_{0},\eta_{1},...,\eta_{m}$ represent the discount factors across various time-scales, and define $D_{0}(s,a):=Q_{\eta_{0}}(s,a)$. The action-value function $Q_{\eta_{m}}(s,a)$ is defined as the cumulative sum of all D-components up to m:
$Q_{\eta_{m}}(s,a)=\sum_{x=0}^{m}D_{x}(s,a)$ (10)
@@
$Q(s_{n},a_{n})\leftarrow Q(s_{n},a_{n})+\alpha\left[r_{n}+\eta Q(s_{n+1},a_{n+1})-Q(s_{n},a_{n})\right]$ (11)
In this context, $r_{n}$ signifies the reward acquired from performing action $a_{n}$ in state $s_{n}$, whereas $(s_{n+1},a_{n+1})$ denotes the ensuing state-action pair. The SARSA update rule is implemented independently to each delta component $D_{n}(s,a)$. The modifications for each component utilize the same architecture as the conventional SARSA approach, but it is tailored to include the delta function at each distinct time-scale. The update in Eq.11 for single-step TD SARSA ($\Delta$) can be articulated with numerous time-scales by employing the principle of splitting the update into several discount factors[8] as demonstrated below:
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}D_{m}(s_{n+1},a_{n+1})\right]$ (12)
where $Q_{\eta_{m}}(s_{n},a_{n})$ represents the action-value function with a discount factor $\eta_{m}$. $D_{m}(s_{n},a_{n})$ denotes the delta function for SARSA ($\Delta$), while $\eta_{m}$ represents the discount factor for time-scale m, while $Q_{\eta_{m}}(s_{n+1},a_{n+1})$ is the SARSA update for action-value associated with the subsequent state-action pair. The single-step TD SARSA ($\Delta$) mirrors the original TD update, but it is applied to action-value functions Q rather than state-value functions V. The single-step Bellman Eq. for the Q-value function in standard SARSA is as follows:
$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\left[r_{n}+\eta_{n}Q_{\eta_{n}}(s_{n+1},a_{n+1})\right]$ (13)
This denotes the expected value of the reward acquired at time n, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{n+1},a_{n+1})$. Currently, rather than utilizing a singular discount factor $\eta$, we are employing a sequence of discount factors $\eta_{0},\eta_{1},.....,\eta_{m}$. The delta function $D_{m}(s_{n},a_{n})$ is defined as the difference between action-value functions associated with successive discount factors from Eq. 9:
$D_{m}(s_{n},a_{n})=Q_{\eta_{m}}(s_{n},a_{n})-Q_{\eta_{m-1}}(s_{n},a_{n})$ (14)
To expand the delta function $D_{z}$ utilizing its definition, we replace the Bellman Eq. for both $Q_{\eta_{m}}$ and $Q_{\eta_{m-1}}$, thereby expressing $D_{m}$ in relation to these action-value functions. Commence by composing:
@@
Since the reward terms $r_{n}$ are included in both formulations, simplify by eliminating them.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}Q_{\eta_{m}}(s_{n+1},a_{n+1})-\eta_{m-1}Q_{\eta_{m-1}}(s_{n+1},a_{n+1})\right]$
@@
$Q_{\eta_{m}}(s_{n+1},a_{n+1})=D_{m}(s_{n+1},a_{n+1})+Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$
Substitute the definition of $D_{m}$ into the Eq. below.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}(D_{m}(s_{n+1},a_{n+1})+Q_{\eta_{m-1}}(s_{n+1},a_{n+1}))-\eta_{n-1}Q_{\eta_{n-1}}(s_{n+1},a_{n+1})\right]$
@@
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}D_{m}(s_{n+1},a_{n+1})+(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})\right]$
Integrate the terms to derive the final update Eq. for $D_{m}(s_{n},a_{n})$.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}D_{m}(s_{n+1},a_{n+1})\right]$ (16)
This Eq. demonstrates that the update for $D_{m}$ integrates the disparity in discount factors $\eta_{m}-\eta_{m-1}$ with respect to the action-value function $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ and the bootstrapping from the subsequent step $D_{m}(s_{n+1},a_{n+1})$. This denotes a Bellman Eq. for $D_{m}$, combining a decay factor $\eta_{m}$ and the reward $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$. Consequently, it can be utilized to define the expected TD update for $D_{m}$. In this expression, $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ can be expressed as the summation of $D_{m}(s_{n+1},a_{n+1})$ for $m\leq M-1$, indicating that the Bellman Eq. for $D_{m}$ is contingent upon the values of all delta functions $D_{m}$ for $m\leq M-1$.
This approach treats the delta value function at each time-scale as an autonomous RL issue, with rewards obtained from the action-value function of the immediately preceding time-scale. Thus, for a target discounted action-value function $Q_{\eta_{m}}(s,a)$, all delta components can be trained simultaneously through a TD update, utilizing the prior values of each estimator for bootstrapping. This process necessitates the assumption of a sequence of discount factors, denoted as $\eta_{m}$, which includes both the minimum and maximum values, $\eta_{0}$ and $\eta_{m}$[8].
@@ 4.3 Multi-Step TD ( SARSA($\Delta$) )
Numerous studies indicate that multi-step TD methods typically exhibit greater efficiency compared to single-step TD methods [29]. In multi-step TD SARSA($\Delta$), rewards are aggregated over multiple steps instead of depending on a single future reward. This approach takes into account the differences between consecutive discount factors, $\eta_{m}$ and $\eta_{m-1}$, while utilizing the value function at state $s_{n+k_{m}}$ for bootstrapping. In SARSA($\Delta$), the update Eqs. are modified to aggregate rewards across a series of transitions prior to bootstrapping from either the Q-values or delta estimators. In multi-step SARSA($\Delta$), the agent aggregates rewards over K steps, leading to a corresponding modification in the update rule for each $D_{m}$.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=1}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{x}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (17)
@@
In this case, $r_{m+x}$ represents the reward obtained at time step $n+x$. $a_{n+k_{m}}$ and the action is chosen greedily as $a=\arg\max_{a}Q(s_{n+k_{m}},a)$, where $k_{m}$ is the number of steps linked to the discount factor $\eta_{m}$. This method generalizes Q-learning to multi-step TD learning by aggregating rewards across several stages and utilizing bootstrapping from both the current and prior time-scale action-value functions. In normal multi-step TD learning, the TD error is computed by aggregating rewards over several stages rather than relying on a single future step, followed by bootstrapping from the action-value at the last step. For example, utilizing a singular discount factor $\eta$, the multi-step TD update equation is expressed as:
$Q(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k-1}\eta^{x}r_{n+x}+\eta^{k}Q(s_{n+k},a_{n+k})\right]$ (18)
@@
Herein, The initial segment of the Eq. aggregates the rewards across k steps, each diminished by $\eta$. The second component derives from the action-value function at time step n+k. We expand this methodology to include various discount factors and delta decomposition, concentrating on representing the delta function $D_{m}(s_{n},a_{n})$ as the difference between action-value functions with discount factors $\eta_{m}$ and $\eta_{m-1}$. According to the definition of $D_{m}$, we obtain:
$D_{m}(s_{n},a_{n})=Q_{\eta_{m}}(s_{n},a_{n})-Q_{\eta_{m-1}}(s_{n},a_{n})$
@@
We now implement the multi-step variant of the Bellman Eq. for both $Q_{\eta_{m}}(s_{n},a_{n})$ and $Q_{\eta_{m-1}}(s_{n},a_{n})$. The multi-step Bellman Eq. for $Q_{\eta_{m}}(s_{n},a_{n})$ is expressed as follows:
$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
For $Q_{\eta_{m-1}}(s_{n},a_{n})$, the multi-step Bellman Eq. is expressed as follows:
$Q_{\eta_{m-1}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
@@
$D_{m}(s_{n},a_{n})=\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]-\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
We will now broaden and simplify the expressions. The immediate reward terms $r_{n+x}$ are present in both Eqs. but are adjusted by distinct discount factors, $\eta_{m}$ and $\eta_{m-1}$. This enables us to express the disparity in rewards as:
$\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}-\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}=\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}$
@@
$\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})-\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})=(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$
@@
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (19)
This Eq. extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL. As a result, each $D_{m}$ obtains a share of the rewards from the environment up to time-step $k_{m}-1$. Furthermore, each $D_{m}$ utilizes its distinct action-value function in conjunction with the value from the prior time-scale. A variation of this algorithm, utilizing k-step bootstrapping as described in [29], is presented in Algorithm 1. Despite Algorithm 1 exhibiting quadratic complexity in relation to M, it can be executed with linear complexity for substantial M by preserving $\hat{Q}$ action-values at each time-scale $\eta_{m}$.
@@
Inputs: Pick out the discount factors ($\eta_{0},\eta_{1},...,\eta_{M}$), bootstrapping steps ($k_{0},k_{1},...,k_{M}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{M}$).
Set the initial value of $D_{m}(s,a)=0$ for all states, actions, and scales m.
@@
     for n = 0, 1, 2,… do $\triangleright$ Loop for each time step
         Take action $a_{n}$, observe reward $r_{n}$ and next state $s_{n+1}$.
         Select action $a_{n+1}$ based on a policy.
         for m = 0, 1, …, M do
              if m = 0 then
                  $G^{0}=\sum_{x=0}^{k_{0}-1}\eta_{0}^{x}r_{n+x}+\eta_{0}^{k_{0}}D_{0}(s_{n+k_{0}},a_{n+k_{0}})$
              $\triangleright$ Utilizing Eq.10, we substitute $Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$ by summing the D-components up to $D_{m-1}$ in Eq.4.3.
                  $G^{m}=\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\sum_{m=0}^{m-1}D_{m-1}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$